Consider the wooden board hinged as shown with the 50 lbf force applied as shown. 50j (lb) .T 4,0,3) ft (a) What is the torque about the x-axis in ft-lbf? About the y-axis? And the z-axis?
Answer #1

X = 4ft , y= 0ft , z = 3ft and Force F = 50j^

Now a.) Torque about z-axis

Tx = x×F = 4i^ × 50j^ = 200 k^ ft-lbf

b.)

Ty = 0×F = 0 ft-lbf

C.) Tz = z × F = 3k^×50j^ = -150i^ ft-lbf